OpenStudy (anonymous):

Find the volume of a sphere that has a surface area of 16 sq. in. cubic inches

OpenStudy (anonymous):

\[4*22/7*r^2 = 16\] \[r^2 = (16*7)/(22*4)\]

OpenStudy (anonymous):

9856

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

\[r^2 = 112/88\]

OpenStudy (anonymous):

1.2?

OpenStudy (anonymous):

\[r = \sqrt{1.27}\]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

\[r = 1.13\]

OpenStudy (anonymous):

\[V = 4/3 \pi r ^3\]

OpenStudy (anonymous):

\[(4/3)*(22/7)*(1.13)^3\]

OpenStudy (anonymous):

V = 6.05 (approx)